### v4

# Mykola Khrypchenko

Abstract
We describe $\sigma$-matching, interchangeable and, as a consequence, totally compatible products on some classes of associative algebras, including unital algebras, the semigroup algebras of rectangular bands, algebras with enough idempotents, free non-unital associative algebras and free non-unital commutative associative algebras.
keywords:
Associative algebras, Compatible structures, $\sigma$-matching structures, Interchangeable structures, Totally compatible structures.
[…]
Introduction
Let $n$ be a positive integer and $\mathcal{A}$ a variety of algebras with one $n$-linear operation $\cdot$$\dots,\cdot$]. Consider two $n$-linear operations $\cdot,\dots,\cdot]_{1}$ and $\cdot,\dots,\cdot]_{2}$ on the same vector space $V$, such that the algebras $(V,[\cdot,\dots,\cdot]_{1})$ and $(V,[\cdot,\dots,\cdot]_{2})$ belong to $\mathcal{A}$. Then $\cdot,\dots,\cdot]_{1}$ and $\cdot,\dots,\cdot]_{2}$ are said to be compatible if $(V,[\cdot,\dots,\cdot]_{1}+[\cdot,\dots,\cdot]_{2})$ belongs to $\mathcal{A}$. In this case, the triple $(V,[\cdot,\dots,\cdot]_{1},[\cdot,\dots,\cdot]_{2})$ is called a compatible $\mathcal{A}$-algebra. For example, there are compatible associative algebras [6, 17], compatible Lie algebras [12], compatible pre-Lie algebras [1], compatible Leibniz algebras [16] and so on. If the signature of a variety $\mathcal{A}$ has several operations (of possibly different types), one can consider the compatibility with respect to some fixed operation, for example, compatible Poisson brackets [3], compatible Hom-Lie algebras [8], compatible Hom-Lie triple systems [21] and so on, or with respect to all the operations at the same time as in [20, Definition A].
The notion of a pair of compatible Poisson or Lie brackets seems to have firstly appeared in mathematical physics [15, 19, 3] and has been studied in this context for a couple of decades [6, 4, 12, 13]. A pair of compatible associative products was introduced (under the name “quantum bi-Hamiltonian system”) and investigated from the purely algebraic point of view in [6], where some general examples of such products were given. Odesski and Sokolov observed in [17] that the description of pairs of compatible associative structures can be an interesting mathematical problem on its own. Thus, they managed to characterize in [17] the compatible associative products on the matrix algebra $M_{n}(\mathbb{C})$ in terms of the so-called $n$-dimensional representations of $M$-algebras, and in the subsequent paper [18] they used these products to construct a solution to the classical Yang-Baxter equation. Recently, there has appeared an interest in the algebraic classification of nilpotent compatible algebras of small dimensions [1, 14].
At a more abstract level, compatible operations have been studied by specialists in operad theory. Dotsenko and Khoroshkin calculated in [10] the dimensions of the graded components of the operads of a pair of compatible Lie brackets and compatible Poisson brackets. For the operad of a pair of compatible associative products, the corresponding dimensions have been calculated by Dotsenko in [9]. Strohmayer [20] gave a general way to construct the (binary quadratic) operad of two compatible structures from the (binary quadratic) operad of a single structure. He also pointed out in [20, p. 2525] three other kinds of compatibilities, the first of which corresponds to the interchangeability of the structures, the second one is a part of the matching compatibility and the third one is the total compatibility. Algebras with two matching associative products (called matching dialgebras) and their connection to other classes of algebras were studied in [26], while totally compatible associative products and the corresponding operads were investigated in [27, 25]. The notions of compatibility, matching compatibility and total compatibility were generalized to families of specific algebras (including the associative ones) in [28] and to families of algebras over an arbitrary (unary binary quadratic/cubic) operad in [24]. Finally, even more general matching compatibilities were introduced and studied in [23].
Observe that, given an associative algebra $(A,\cdot)$, the compatible (with $\cdot$) products on $A$ are exactly the associative Hochschild $2$-cocycles of $A$ with values in $A$. However, even knowing $HH^{2}(A,A)$, it is difficult to characterize the associativity of the corresponding $2$-cocycles in simple terms (for example, see the case $HH^{2}(A,A)={0}$ in [17]). Nevertheless, the description of matching and interchangeable products on $A$ seems to be a manageable task. Moreover, it turns out that for unital associative algebras such structures admit simple descriptions. In the non-unital case there are still interesting algebras to consider.
In Section˜1 we specify the definitions of $\sigma$-matching ($\sigma\in S_{2}$), interchangeable and total compatibility to the case of a pair of associative products and give examples showing the difference between the compatibilities. As often happens, nilpotent algebras are a good source of such examples. The case of unital products is considered in Section˜2 (see Propositions˜2.2 and 2.6) as a motivation for Section˜3, where we prove some general results on idempotent algebras (see Propositions˜3.1 and 3.2) and proceed to more specific classes of such algebras: the semigroup algebras of rectangular bands and algebras with enough idempotents (the latter class includes some algebras of infinite matrices, path algebras and infinite direct sums of unital algebras). The descriptions of $\sigma$-matching and interchangeable products on these algebras are given in Propositions˜3.3, 3.8, 3.18 and 3.23. Section˜4 is devoted to free non-unital associative algebras. We first prove a general result valid for all algebras without zero divisors (Proposition˜4.1) and then treat separately the cases of non-commutative and commutative non-unital free algebras (see Propositions˜4.2, 4.6, 4.9 and 4.12).
1 Definitions and preliminaries
All the algebras and vector spaces will be over a field $K$ and all the products will be binary and bilinear.
1.1 Compatible structures
Given two bilinear binary operations $\cdot_{1}$ and $\cdot_{2}$ on a vector space $V$, their sum $\star=\cdot_{1}+\cdot_{2}$ is defined by
$a\star b=a\cdot_{1}b+a\cdot_{2}b.$
Recall that two associative products $\cdot_{1}$ and $\cdot_{2}$ on $V$ are called compatible, if $\cdot_{1}+\cdot_{2}$ is also associative. This is equivalent to the following equality:
$(a\cdot_{1}b)\cdot_{2}c+(a\cdot_{2}b)\cdot_{1}c=a\cdot_{1}(b\cdot_{2}c)+a\cdot_{2}(b\cdot_{1}c)$ (1)
for all $a,b,c\in V$. As it was observed in [6, Remark on p. 4801] and can be seen directly, $\cdot_{1}$ and $\cdot_{2}$ are compatible if and only if $\cdot_{1}$ is a Hochschild $2$-cocycle of $(V,\cdot_{2})$ with values in $V$ (or, symmetrically, $\cdot_{2}$ is a Hochschild $2$-cocycle of $(V,\cdot_{1})$ with values in $V$).
Let us consider the following particular cases of ˜1.
Definition 1.1.
Adopting the terminology of [23] we say that two associative products $\cdot_{1}$ and $\cdot_{2}$ on $V$ are
  1. (i)
    $\sigma$-matching (where $\sigma\in S_{2}=\{\mathrm{id},(12)\}$), if
[…]
  2. (ii)
    totally compatible, if
    $(a\cdot_{1}b)\cdot_{2}c=(a\cdot_{2}b)\cdot_{1}c=a\cdot_{1}(b\cdot_{2}c)=a\cdot_{2}(b\cdot_{1}c),$ (3)
for all $a,b,c\in V$.
Usually $\mathrm{id}$-matching $\cdot_{1}$ and $\cdot_{2}$ are simply called matching in the literature. Observe that in this case $(V,\cdot_{1},\cdot_{2})$ is a matching dialgebra [26], also known as $As^{(2)}$-algebra [29]. We will consider one more “compatibility-type” condition which makes sense for not necessarily associative products.
Definition 1.2.
We say that two (not necessarily associative) products $\cdot_{1}$ and $\cdot_{2}$ on $V$ are interchangeable if
[…]
Remark 1.3.
For two associative products $\cdot_{1}$ and $\cdot_{2}$ on $V$ the following are equivalent:
  1. (i)
    $\cdot_{1}$ and $\cdot_{2}$ are totally compatible;
  2. (ii)
    $\cdot_{1}$ and $\cdot_{2}$ are interchangeable and at least one of the two equalities ˜2 holds for some $\sigma\in S_{2}$;
  3. (iii)
    $\cdot_{1}$ and $\cdot_{2}$ are $\sigma$-matching for some $\sigma\in S_{2}$ and at least one of the two equalities ˜4 holds;
  4. (iv)
    $\cdot_{1}$ and $\cdot_{2}$ are $\sigma_{1}$-matching for some $\sigma_{1}\in S_{2}$ and at least one of the two equalities ˜2 holds for $\sigma\neq\sigma_{1}$;
If $\mathrm{char}(K)\neq 2$, then each of the conditions ˜(i), (ii), (iii) and (iv) is also equivalent to
  1. (v)
    $\cdot_{1}$ and $\cdot_{2}$ are interchangeable and compatible.
Let $(A,\cdot)$ be an associative algebra. By a compatible (resp. $\sigma$-matching, interchangeable or totally compatible) structure on $A$ we mean an associative product * on $A$ that is compatible (resp. $\sigma$-matching, interchangeable or totally compatible) with $\cdot$.
The following examples show that the classes of $\mathrm{id}$-matching, $(12)$-matching and interchangeable structures on $(A,\cdot)$ are different and no class is in general contained in another one (although all of them contain the totally compatible structures).
Example 1.4.
Let $A$ be the $3$-dimensional nilpotent associative algebra with a basis $\{e_{i}\}_{i=1}^{3}$ and the multiplication table $e_{1}\cdot e_{2}=e_{3}$.
The product * on $A$ given by $e_{1}*e_{1}=e_{1}$ and $e_{1}*e_{2}=e_{2}$ is clearly associative (observe that (A,)(A,*) is isomorphic to the direct sum of a $1$-dimensional algebra with zero multiplication and the subalgebra of $M_{2}(K)$ generated by the matrix units $\{E_{11},E_{12}\}$). Then $(e_{i}*e_{j})\cdot e_{k}\neq 0\Leftrightarrow e_{i}\cdot(e_{j}*e_{k})\neq 0\Leftrightarrow(i,j,k)=(1,1,2)$, in which case $(e_{i}*e_{j})\cdot e_{k}=e_{i}\cdot(e_{j}*e_{k})=e_{3}$. Moreover, $(A\cdot A)*A=A*(A\cdot A)=\{0\}$. Thus, the product * is $(12)$-matching with $\cdot$, but it is neither $\mathrm{id}$-matching nor interchangeable with $\cdot$ because $(A*A)\cdot A\neq\{0\}$.
On the other hand, consider the associative product $\star$ on $A$ given by $e_{1}\star e_{1}=e_{1}$ and $e_{1}\star e_{3}=e_{3}$. Then,
$(e_{i}\star e_{j})\cdot e_{k}\neq 0\quad\textup{if and only if}\quad e_{i}\star(e_{j}\cdot e_{k})\neq 0\quad\textup{if and only if}\quad(i,j,k)=(1,1,2),$
in which case $(e_{i}\star e_{j})\cdot e_{k}=e_{i}\star(e_{j}\cdot e_{k})=e_{3}$. Moreover, $(A\cdot A)\star A=A\cdot(A\star A)=\{0\}$. Thus, the product $\star$ is $\mathrm{id}$-matching with $\cdot$, but it is neither $(12)$-matching nor interchangeable with $\cdot$ because $(A\star A)\cdot A\neq\{0\}$.
Example 1.5.
Let $A$ be the algebra with a basis $\{e_{i}\}_{i=1}^{6}$ and multiplication table
$e_{1}\cdot e_{2}=e_{4},\>e_{1}\cdot e_{5}=e_{6},\>e_{4}\cdot e_{3}=e_{6},\>e_{2}\cdot e_{3}=e_{5}.$
It is a nilpotent associative algebra isomorphic to the subalgebra of $M_{4}(K)$ generated by the matrix units $\{E_{12},E_{23},E_{34}\}$.
The product * on $A$ given by $e_{1}*e_{2}=e_{5}$, $e_{1}*e_{4}=e_{6}$ is clearly associative because $(A*A)*A=A*(A*A)=\{0\}$. Furthermore,
$e_{i}*(e_{j}\cdot e_{k})\neq 0\Leftrightarrow e_{i}\cdot(e_{j}*e_{k})\neq 0\Leftrightarrow(i,j,k)=(1,1,2),$
in which case $e_{i}*(e_{j}\cdot e_{k})=e_{i}\cdot(e_{j}*e_{k})=e_{6}$. Moreover, $(A*A)\cdot A=(A\cdot A)*A=\{0\}$. Thus, * is interchangeable with $\cdot$, but it is neither $\mathrm{id}$-matching nor $(12)$-matching with $\cdot$ because $A*(A\cdot A)\neq\{0\}$.
Given two structures $_{1}$ and $_{2}$ on $A$ that are compatible with $\cdot$, we say that $_{1}$ and $_{2}$ are isomorphic, if there exists an automorphism $\phi$ of $(A,\cdot)$ such that
$\phi(a*_{1}b)=\phi(a)*_{2}\phi(b)$
for all $a,b\in A$ (i.e. if $(A,\cdot,*_{1})$ and $(A,\cdot,*_{2})$ are isomorphic).
2 $\sigma$-matching and interchangeable structures on unital associative algebras
Although unital associative algebras are a subclass of algebras with enough idempotents considered in Section˜3, we begin with this particular case to motivate the choice of the context for Section˜3. We also introduce here some notions that will be used below.
For any associative algebra $(A,\cdot)$ and fixed $x\in A$, one defines the mutation [11] of $\cdot$ by $x$ to be the product $\cdot_{x}$ on $A$ given by $a\cdot_{x}b=a\cdot x\cdot b$ for all $a,b\in A$. The following should be well-known (cf. [6, Formula (9)] and [17, Example 2.2]), but we couldn’t find an explicit proof.
Lemma 2.1.
Let $(A,\cdot)$ be an associative algebra. For any $x\in A$ the product $\cdot_{x}$ is associative and $\mathrm{id}$-matching with $\cdot$.
Proof.
For $a,b,c\in A$ we have $(a\cdot_{x}b)\cdot c=a\cdot x\cdot b\cdot c=a\cdot_{x}(b\cdot c)$ and $(a\cdot b)\cdot_{x}c=a\cdot b\cdot x\cdot c=a\cdot(b\cdot_{x}c)$. Moreover, $(a\cdot_{x}b)\cdot_{x}c=a\cdot x\cdot b\cdot x\cdot c=a\cdot_{x}(b\cdot_{x}c)$. ∎
For unital associative algebras the converse also holds.
Proposition 2.2.
Let $(A,\cdot)$ be a unital associative algebra. Then the $\mathrm{id}$-matching structures on $A$ are exactly the mutations of $\cdot$.
Proof.
Let $1$ be the identity element of $(A,\cdot)$ and * an $\mathrm{id}$-matching associative product on $A$. Then, for all $a,b\in A$ we have
$a*b=(a\cdot 1)*b=a\cdot(1*b)=a\cdot(1*(1\cdot b))=a\cdot((1*1)\cdot b)=a\cdot_{1*1}b.\qed$
Let $(A,\cdot)$ be a (not necessarily associative) algebra. Recall that the centroid of $A$ is the space $\Gamma(A)$ of linear maps $\varphi:A\to A$ such that
$x\cdot\varphi(y)=\varphi(x\cdot y)=\varphi(x)\cdot y$
for all $x,y\in A$.
Lemma 2.3.
Let $(A,\cdot)$ be a (not necessarily associative) algebra. Given $\varphi\in\Gamma(A)$, the product
$a*_{\varphi}b:=\varphi(a\cdot b)$ (5)
is interchangeable with $\cdot$. Moreover, if $\cdot$ is associative, then $_{\varphi}$ is also associative and totally compatible with $\cdot$.
Proof.
For all $a,b,c\in A$ we have
$(a*_{\varphi}b)\cdot c$ $=\varphi(a\cdot b)\cdot c=\varphi((a\cdot b)\cdot c)=(a\cdot b)*_{\varphi}c\text{ and }$
$a*_{\varphi}(b\cdot c)$ $=\varphi(a\cdot(b\cdot c))=a\cdot\varphi(b\cdot c)=a\cdot(b*_{\varphi}c).$
The second statement follows from [27, Proposition 2.7]. ∎
Definition 2.4.
The product of the form ˜5 is said to be determined by $\varphi\in\Gamma(A)$.
In the associative case any element $c$ of the center $C(A)$ of $A$ defines $\varphi\in\Gamma(A)$ by means of $\varphi(a)=c\cdot a$. If, moreover, $A$ is unital, then this gives an isomorphism of $K$-spaces $C(A)\cong\Gamma(A)$, whose inverse maps $\varphi\in\Gamma(A)$ to $\varphi(1)$. Thus, we have the following.
Lemma 2.5.
Let $(A,\cdot)$ be a unital associative algebra. Then the products * determined by elements of $\Gamma(A)$ are exactly the mutations of $\cdot$ by elements of $C(A)$.
Proposition 2.6.
Let $(A,\cdot)$ be a unital associative algebra and * be an associative product on $A$. Then the following are equivalent:
  1. (i)
    * is $(12)$-matching with $\cdot$;
  2. (ii)
    * is interchangeable with $\cdot$;
  3. (iii)
    * is totally compatible with $\cdot$;
  4. (iv)
    * is determined by an element of $\Gamma(A)\cong C(A)$.
Proof.
[…]
$1*a=(1*a)\cdot 1=1\cdot(a*1)=a*1.$
It follows that
$(1*1)\cdot a=1\cdot(1*a)=1*a=a*1=(a*1)\cdot 1=a\cdot(1*1),$
so $1*1\in C(A)$. Using these equalities, we also have
$a*b=(a*b)\cdot 1=a\cdot(b*1)=a\cdot((1*1)\cdot b)=a\cdot_{1*1}b$
for all $a,b\in A$.
[…]
Now, $1*a=(1*a)\cdot 1=(1\cdot a)*1=a*1$, whence $(1*1)\cdot a=a\cdot(1*1)$, i.e. $1*1\in C(A)$. Thus,
$a*b=(a\cdot 1)*b=(a*1)\cdot b=a\cdot(1*1)\cdot b=a\cdot_{1*1}b$
for all $a,b\in A$.
[…]
Corollary 2.7.
Let $A$ be a non-commutative unital associative algebra. Then there exist products on $A$ that are $\mathrm{id}$-matching with $\cdot$, but not totally compatible with $\cdot$.
Proof.
Let $x$ be a non-central element of $A$. Then $\cdot_{x}$ is $\mathrm{id}$-matching with $\cdot$ by Lemma˜2.1. If $\cdot_{x}$ was totally compatible with $\cdot$, then by Proposition˜2.6 there would be $c\in C(A)$ such that $\cdot_{x}=\cdot_{c}$, whence $x=1\cdot_{x}1=1\cdot_{c}1=c$, a contradiction. ∎
Remark 2.8.
Let $(A,\cdot)$ be a unital associative algebra and $x,y\in A$. Then the structures $\cdot_{x}$ and $\cdot_{y}$ are isomorphic if and only if there is $\phi\in\operatorname{Aut}(A)$ such that $\phi(x)=y$.
For, given $\phi\in\operatorname{Aut}(A)$, one has
$\phi(a\cdot_{x}b)=\phi(a)\cdot_{y}\phi(b)\Leftrightarrow\phi(a)\cdot\phi(x)\cdot\phi(b)=\phi(a)\cdot y\cdot\phi(b).$
The latter holds for all $a,b\in A$ if and only if $\phi(x)=y$ (take $a=b=1$).
Remark 2.9.
Whenever $\Gamma(A)=K$, there are only two non-isomorphic structures of the form ˜5: the original product $\cdot$ and the zero product.
For, in this case $a*_{\varphi}b=\lambda(a\cdot b)$ for some $\lambda\in K$, so, if $\lambda\neq 0$, the map $\lambda{}^{-1}\mathrm{id}$ is an isomorphism between $\cdot$ and $_{\varphi}$.
Remark 2.10.
For non-unital algebras the result of Proposition˜2.6 may be false. For example, if $A$ is an algebra with zero multiplication, then any associative product * on $A$ is totally compatible with $\cdot$ since all the mixed monomials in ˜3 are zero. However, all the mutations of $\cdot$ and all the products on $A$ determined by elements of $\Gamma(A)$ are zero.
3 $\sigma$-matching and interchangeable structures on certain idempotent associative algebras
In this section we are going to see to which extent the results of Section˜2 generalize to several classes of associative algebras that are in some sense close to being unital. Recall that a (not-necessarily associative) algebra $(A,\cdot)$ is said to be idempotent [^1] In the context of Lie algebras one prefers to use the term “perfect” rather than “idempotent”. if $A\cdot A=A$.
Proposition 3.1.
Let $(A,\cdot)$ be an idempotent associative algebra and * be an associative product on $A$. Then the following are equivalent:
  1. (i)
    * is interchangeable with $\cdot$;
  2. (ii)
    * is totally compatible with $\cdot$.
Proof.
[…]
$a\cdot(b*c)=a\cdot((b_{1}\cdot b_{2})*c)=a\cdot((b_{1}*b_{2})\cdot c)=(a\cdot(b_{1}*b_{2}))\cdot c=(a*(b_{1}\cdot b_{2}))\cdot c=(a*b)\cdot c.$
Since $A\cdot A=A$, then by linearity $a\cdot(b*c)=(a*b)\cdot c$ for all $a,b,c\in A$. Thus, * is totally compatible with $\cdot$ by Remark˜1.3˜(ii). ∎
It is natural to ask if one can replace interchangeable products by $\sigma$-matching ones in Proposition˜3.1. For $\sigma=\mathrm{id}$ the answer is “no” even in the case of unital algebras, as we saw in Corollary˜2.7. For $\sigma=(12)$ there are classes of idempotent algebras containing the unital ones for which the answer is positive and those for which it is negative. We first point out a class admitting the positive answer.
Proposition 3.2.
Let $(A,\cdot)$ be an associative algebra with a left or right unit and * be an associative product on $A$. Then the following are equivalent:
  1. (i)
    * is $(12)$-matching with $\cdot$;
  2. (ii)
    * is totally compatible with $\cdot$.
Proof.
[…]
$(a*b)\cdot c$ $=e\cdot((a*b)\cdot c)$
$=(e\cdot(a*b))\cdot c$
$=((e*a)\cdot b)\cdot c$
$=(e*a)\cdot(b\cdot c)$
$=e\cdot(a*(b\cdot c))$
$=a*(b\cdot c).$
It follows that * is totally compatible with $\cdot$ by Remark˜1.3˜(iv). If $A$ has a right unit $e$, then we symmetrically have
$a\cdot(b*c)$ $=(a\cdot(b*c))\cdot e$
$=a\cdot((b*c)\cdot e)$
$=a\cdot(b\cdot(c*e))$
$=(a\cdot b)\cdot(c*e)$
$=((a\cdot b)*c)\cdot e=(a\cdot b)*c.\qed$
3.1 The semigroup algebra of a rectangular band
Given arbitrary non-empty sets $I$ and $J$, one easily sees that $S:=\{e_{ij}\mid(i,j)\in I\times J\}$ is a semigroup under the multiplication
$e_{ij}\cdot e_{kl}=e_{il}$ (6)
for all $(i,j),(k,l)\in I\times J$. It is called a rectangular band [7], and it is a classical example of a semigroup in which every element is idempotent. Let $(A,\cdot)$ be the semigroup $K$-algebra of $S$. Thus, $A$ is an associative algebra admitting a basis consisting of the idempotents $e_{ij},(i,j)\in I\times J$. We are going to describe the $\sigma$-matching structures on $(A,\cdot)$.
Proposition 3.3.
The $\mathrm{id}$-matching structures on $(A,\cdot)$ are exactly the products * of the form
$e_{ij}*e_{kl}=\lambda_{jk}e_{il},$ (7)
where $\lambda_{jk}\in K$ and $(i,j),(k,l)\in I\times J$.
Proof.
Let * be $\mathrm{id}$-matching with $\cdot$. Then for all $(i,j),(k,l)\in I\times J$ we have
$e_{ij}*e_{kl}=(e_{ij}\cdot e_{ij})*e_{kl}=e_{ij}\cdot(e_{ij}*e_{kl})=e_{ij}\cdot(e_{ij}*(e_{kl}\cdot e_{kl}))=e_{ij}\cdot(e_{ij}*e_{kl})\cdot e_{kl},$
which equals $\mu_{ij,kl}e_{il}$ for some $\mu_{ij,kl}\in K$ by ˜6. Moreover, for any $(p,q)\in I\times J$ we have
$\mu_{ij,kl}e_{iq}=(e_{ij}*e_{kl})\cdot e_{pq}=e_{ij}*(e_{kl}\cdot e_{pq})=e_{ij}*e_{kq}=\mu_{ij,kq}e_{iq},$
whence $\mu_{ij,kl}=\mu_{ij,kq}$. Similarly,
$\mu_{ij,kl}e_{pl}=e_{pq}\cdot(e_{ij}*e_{kl})=(e_{pq}\cdot e_{ij})*e_{kl}=e_{pj}*e_{kl}=\mu_{pj,kl}e_{pl},$
whence $\mu_{ij,kl}=\mu_{pj,kl}$. Thus, $\mu_{ij,kl}$ does not depend on $i$ and $l$, so denoting it by $\lambda_{jk}$, we get ˜7.
Conversely, let * be given by ˜7. Then for all $(i,j),(k,l),(p,q)\in I\times J$ we have
$(e_{ij}*e_{kl})*e_{pq}=\lambda_{jk}e_{il}*e_{pq}=\lambda_{jk}\lambda_{lp}e_{iq}\quad\textup{and}\quad e_{ij}*(e_{kl}*e_{pq})=\lambda_{lp}e_{ij}*e_{kq}=\lambda_{lp}\lambda_{jk}e_{iq},$
so that * is associative. Since
$(e_{ij}*e_{kl})\cdot e_{pq}=\lambda_{jk}e_{il}\cdot e_{pq}=\lambda_{jk}e_{iq}=e_{ij}*e_{kq}=e_{ij}*(e_{kl}\cdot e_{pq})$
and similarly
$(e_{ij}\cdot e_{kl})*e_{pq}=e_{il}*e_{pq}=\lambda_{lp}e_{iq}=e_{ij}\cdot\lambda_{lp}e_{kq}=e_{ij}\cdot(e_{kl}*e_{pq}),$
the product * is $\mathrm{id}$-matching with $\cdot$. ∎
Remark 3.4.
The product ˜7 is a mutation of $\cdot$ if and only if $\lambda_{jk}$ does not depend on $j$ and $k$.
Indeed, given $a=\sum_{p,q}a_{pq}e_{pq}$ with $a_{pq}\in K$, we have $e_{ij}\cdot_{a}e_{kl}=e_{ij}\cdot a\cdot e_{kl}=\lambda e_{il}$, where $\lambda=\sum_{p,q}a_{pq}$. Conversely, if $\lambda_{jk}=\lambda$ for all $j$ and $k$ in ˜7, then $e_{ij}*e_{kl}=\lambda e_{il}=e_{ij}\cdot_{a}\cdot e_{kl}$, where $a=\lambda e_{pq}$ for some fixed arbitrary $(p,q)\in I\times J$.
Lemma 3.5.
Let * be an associative product on $A$. If * is $(12)$-matching with $\cdot$, then
$e_{ij}*e_{kl}=e_{il}*e_{il}$ (8)
for all $(i,j),(k,l)\in I\times J$.
Proof.
We have
$e_{ij}*e_{kl}$ $=(e_{ij}\cdot e_{ij})*e_{kl}$
$=e_{ij}*(e_{ij}\cdot e_{kl})$
$=e_{ij}*e_{il}$
$=e_{ij}*(e_{il}\cdot e_{il})$
$=(e_{ij}\cdot e_{il})*e_{il}$
$=e_{il}*e_{il}.\qed$
Denote by $\operatorname{Ann}(A)$ the (two-sided) annihilator of $A$, i.e. the ideal of $A$ consisting of $a\in A$ such that $a\cdot b=b\cdot a=0$ for all $b\in A$.
Lemma 3.6.
For arbitrary $(i,j)\in I\times J$ the centralizer of $e_{ij}$ in $A$ coincides with $\operatorname{span}_{K}\{e_{ij}\}\oplus\operatorname{Ann}(A)$.
Proof.
It is obvious that $\operatorname{span}_{K}\{e_{ij}\}\cap\operatorname{Ann}(A)=\{0\}$ and that any element of the sum $\operatorname{span}_{K}\{e_{ij}\}+\nolinebreak\operatorname{Ann}(A)$ commutes with $e_{ij}$, so we only need to prove that the centralizer of $e_{ij}$ is contained in $\operatorname{span}_{K}\{e_{ij}\}+\operatorname{Ann}(A)$. Let $a=\sum_{k,l}a_{kl}e_{kl}\in A$ with $a_{kl}\in K$. Then $a\cdot e_{ij}=\sum_{k,l}a_{kl}e_{kj}=\sum_{k\in I}\left(\sum_{l\in J}a_{kl}\right)e_{kj}$ and $e_{ij}\cdot a=\sum_{k,l}a_{kl}e_{il}=\sum_{l\in J}\left(\sum_{k\in I}a_{kl}\right)e_{il}$. Thus,
$a\cdot e_{ij}=e_{ij}\cdot a\Leftrightarrow\begin{cases}\sum_{l\in J}a_{kl}=0\text{ for all }k\neq i,\\ \sum_{k\in I}a_{kl}=0\text{ for all }l\neq j,\\ \sum_{l\in J}a_{il}=\sum_{k\in I}a_{kj}.\end{cases}$ (9)
Assuming that $a\cdot e_{ij}=e_{ij}\cdot a$, for all $(p,q)\in I\times J$ by ˜9 we have
$a\cdot e_{pq}=\sum_{k\in I}\left(\sum_{l\in J}a_{kl}\right)e_{kq}=\lambda e_{iq}$
and
$e_{pq}\cdot a=\sum_{l\in J}\left(\sum_{k\in I}a_{kl}\right)e_{pl}=\lambda e_{pj},$
where $\lambda=\sum_{l\in J}a_{il}=\sum_{k\in I}a_{kj}$. Then $(a-\lambda e_{ij})\cdot e_{pq}=a\cdot e_{pq}-\lambda e_{iq}=0$ and $e_{pq}\cdot(a-\lambda e_{ij})=e_{pq}\cdot a-\lambda e_{pj}=0$, so that $a-\lambda e_{ij}\in\operatorname{Ann}(A)$. ∎
Lemma 3.7.
Let * be an associative product on $A$. If * is $(12)$-matching with $\cdot$, then there exist $\lambda\in K$ and $r_{ij}\in\operatorname{Ann}(A)$, such that
$e_{ij}*e_{ij}=\lambda e_{ij}+r_{ij}$ (10)
for all $(i,j)\in I\times J$.
Proof.
Since $e_{ij}\cdot(e_{ij}*e_{ij})=(e_{ij}*e_{ij})\cdot e_{ij}$, by Lemma˜3.6 there are $\lambda_{ij}\in K$ and $r_{ij}\in\operatorname{Ann}(A)$, such that $e_{ij}*e_{ij}=\lambda_{ij}e_{ij}+r_{ij}$. Now, on the one hand $(e_{ij}*e_{ij})\cdot e_{il}=(\lambda_{ij}e_{ij}+r_{ij})\cdot e_{il}=\lambda_{ij}e_{il}$, and on the other hand using ˜8 we have
$(e_{ij}*e_{ij})\cdot e_{il}=e_{ij}\cdot(e_{ij}*e_{il})=e_{ij}\cdot(e_{il}*e_{il})=e_{ij}\cdot(\lambda_{il}e_{il}+r_{il})=\lambda_{il}e_{il},$
whence $\lambda_{ij}=\lambda_{il}$ for all $i\in I$ and $j,l\in J$. Similarly, $e_{kj}\cdot(e_{ij}*e_{ij})=e_{kj}\cdot(\lambda_{ij}e_{ij}+r_{ij})=\lambda_{ij}e_{kj}$ and
$e_{kj}\cdot(e_{ij}*e_{ij})=(e_{kj}*e_{ij})\cdot e_{ij}=(e_{kj}*e_{kj})\cdot e_{ij}=(\lambda_{kj}e_{kj}+r_{kj})\cdot e_{ij}=\lambda_{kj}e_{kj}$
imply $\lambda_{ij}=\lambda_{kj}$ for all $i,k\in I$ and $j\in J$. Thus, $\lambda_{ij}=\lambda_{kl}$ for all $(i,j),(k,l)\in I\times J$, and ˜10 follows. ∎
Proposition 3.8.
A bilinear product * on $A$ is a $(12)$-matching structure on $(A,\cdot)$ if and only if * is associative and, for all $(i,j),(k,l)\in I\times J$,
$e_{ij}*e_{kl}=\lambda e_{il}+r_{il},$ (11)
where $\lambda\in K$ and $r_{il}\in\operatorname{Ann}(A)$.
Proof.
By Lemmas˜3.5 and 3.7 any $(12)$-matching structure on $(A,\cdot)$ has the form ˜11. Conversely, let * be a product on $A$ given by ˜11. Then for all $(i,j),(k,l),(p,q)\in I\times J$ we have $(e_{ij}*e_{kl})\cdot e_{pq}=(\lambda e_{il}+r_{il})\cdot e_{pq}=\lambda e_{iq}$ and $e_{ij}\cdot(e_{kl}*e_{pq})=e_{ij}\cdot(\lambda e_{kq}+r_{kq})=\lambda e_{iq}$. Moreover, $(e_{ij}\cdot e_{kl})*e_{pq}=e_{il}*e_{pq}=\lambda e_{iq}+r_{iq}$ and $e_{ij}*(e_{kl}\cdot e_{pq})=e_{ij}*e_{kq}=\lambda e_{iq}+r_{iq}$. ∎
Corollary 3.9.
The totally compatible structures on $(A,\cdot)$ are exactly the products * of the form
$e_{ij}*e_{kl}=\lambda e_{il},$ (12)
where $\lambda\in K$ and $(i,j),(k,l)\in I\times J$.
Proof.
Let * be totally compatible with $\cdot$. In particular, * is $(12)$-matching with $\cdot$. Then for all $(i,j)\in I\times J$ by ˜11 we have $(e_{ij}*e_{ij})\cdot e_{ij}=\lambda e_{ij}$, while $e_{ij}*(e_{ij}\cdot e_{ij})=e_{ij}*e_{ij}=\lambda e_{ij}+r_{ij}$. Thus, $r_{ij}=0$ for all $(i,j)\in I\times J$ in ˜11, and we get ˜12. Conversely, the product ˜12 is $_{\varphi}$, where $\varphi=\lambda\mathrm{id}\in\Gamma(A)$, so * is totally compatible with $\cdot$ by Lemma˜2.3. ∎
Corollary 3.10.
We have $\Gamma(A)=K$.
Proof.
Let $\varphi\in\Gamma(A)$ and consider the totally compatible product $_{\varphi}$ on $A$ as in Lemma˜2.3. By Corollary˜3.9 there exists $\lambda\in K$ such that $\varphi(e_{ij})=\varphi(e_{ij}\cdot e_{ij})=e_{ij}*_{\varphi}e_{ij}=\lambda e_{ij}$ for all $(i,j)\in I\times J$. By linearity, $\varphi=\lambda\mathrm{id}$. ∎
We summarize the results on totally compatible structures on $A$ in the following.
Proposition 3.11.
Let * be an associative product on $A$. Then the following are equivalent:
  1. (i)
    * is interchangeable with $\cdot$;
  2. (ii)
    * is totally compatible with $\cdot$;
  3. (iii)
    * is determined by some $\varphi\in\Gamma(A)$.
Proof.
This is a consequence of Propositions˜3.1, 2.3 and 3.9. ∎
In general, the annihilator part of ˜11 may be non-trivial, which makes it difficult to characterize explicitly the associativity of ˜11 in terms of $\lambda$ and $r_{il}$. So, we just give an example of such an associative product below, which also provides a $(12)$-matching structure on $A$ that is not totally compatible with $\cdot$, showing that one cannot replace interchangeable structures by $(12)$-matching ones in Proposition˜3.1.
Example 3.12.
Let $I=J=\{1,2\}$ and $A$ be the semigroup algebra of the rectangular band $I\times J$. Then $\operatorname{Ann}(A)=\operatorname{span}_{K}\{e_{11}-e_{12}-e_{21}+e_{22}\}$. Indeed, if $a=\sum_{i,j}a_{ij}e_{ij}$ with $a_{ij}\in K$ belongs to $\operatorname{Ann}(A)$, then $0=a\cdot e_{11}=(a_{11}+a_{12})e_{11}+(a_{21}+a_{22})e_{21}$ and $0=e_{11}\cdot a=(a_{11}+a_{21})e_{11}+(a_{12}+a_{22})e_{12}$, whence $a_{11}=-a_{12}=-a_{21}=a_{22}$, so that $a\in\operatorname{span}_{K}\{e_{11}-e_{12}-e_{21}+e_{22}\}$. Conversely, $e_{11}-e_{12}-e_{21}+e_{22}\in\operatorname{Ann}(A)$ by a straightforward calculation.
Consider the product * on $A$ given by
$e_{ij}*e_{kl}=\begin{cases}e_{11}-e_{12}-e_{21}+e_{22},&(i,l)=(1,2),\\ 0,&\text{otherwise}.\end{cases}$
Then $(A*A)*A=A*(A*A)=\{0\}$, so that (A,)(A,*) is associative. Observe that * is of the form ˜11 for $\lambda=0$ and $r_{il}=e_{11}-e_{12}-e_{21}+e_{22}\in\operatorname{Ann}(A)$ for $(i,l)=(1,2)$ and $r_{il}=0$ otherwise. Thus, * is $(12)$-matching with $\cdot$, but it is not totally compatible with $\cdot$ by Corollary˜3.9.
However, in some cases $\operatorname{Ann}(A)=\{0\}$ and we have the following result.
Proposition 3.13.
Let $|I|=1$ or $|J|=1$ (so that $S$ is a right zero or a left zero semigroup, respectively). For an associative product * on $A$ the following are equivalent:
  1. (i)
    * is $(12)$-matching with $\cdot$;
  2. (ii)
    * is interchangeable with $\cdot$;
  3. (iii)
    * is totally compatible with $\cdot$;
  4. (iv)
    * is determined by some $\varphi\in\Gamma(A)$.
Proof.
[…]
3.2 Algebras with enough idempotents
An associative algebra $A$ is said to have enough idempotents [22] if there is a family $E$ of orthogonal idempotents $\{e_{i}\}_{i\in I}\subseteq A$ such that $A=\bigoplus_{i\in I}e_{i}A=\bigoplus_{i\in I}Ae_{i}$. It follows that $A=\bigoplus_{i,j\in I}e_{i}Ae_{j}$. Observe that $A$ is an idempotent algebra, and it is unital if and only if $|I|<\infty$, in which case $\sum_{i\in I}e_{i}$ is the unit of $A$.
Example 3.14.
The following non-unital associative $K$-algebras have enough idempotents.
  1. (i)
    The algebra of infinite matrices over $K$ with a finite number of nonzero entries. The corresponding idempotents are the matrix units $E_{ii}$.
  2. (ii)
    The path algebra of a quiver with infinite number of vertices. The corresponding idempotents are the trivial paths $e_{x}$, where $x$ is a vertex.
  3. (iii)
    The direct sum of an infinite family of unital algebras. The corresponding idempotents are the units of the direct summands. Any algebra with enough idempotents such that $E\subseteq C(A)$ is of this form.
Let $(A,\cdot)$ be an algebra with enough idempotents. Given $a\in A$, we write $a=\sum_{i,j\in I}a_{ij}$, where $a_{ij}=e_{i}ae_{j}\in e_{i}Ae_{j}$. Define $M$ to be the $K$-space $\prod_{i,j\in I}e_{i}Ae_{j}$ whose elements will be denoted by $m=\prod_{i,j\in I}m_{ij}$ with $m_{ij}\in e_{i}Ae_{j}$. We will identify $A$ with the subspace of $M$ consisting of $m\in M$ with finite $\operatorname{supp}(m):=\{(i,j)\in I^{2}\mid m_{ij}\neq 0\}$.
Proposition 3.15.
The algebra structure on $A$ extends to an $A$-bimodule structure on $M$ by means of
[…]
for all $a\in A$ and $m\in M$. Moreover, $AMA=A$.
Proof.
The operations ˜13 are clearly well-defined and bilinear. For all $a,b\in A$ and $m\in M$ we have
$(a(bm))_{ij}$ $=\sum_{k\in I}a_{ik}(bm)_{kj}=\sum_{k\in I}\sum_{l\in I}a_{ik}b_{kl}m_{lj}$
$=\sum_{l\in I}\sum_{k\in I}a_{ik}b_{kl}m_{lj}=\sum_{l\in I}(ab)_{il}m_{lj}=((ab)m)_{ij},$
and similarly one proves that $((ma)b)_{ij}=(m(ab))_{ij}$. Finally,
$(a(mb))_{ij}=\sum_{k\in I}\sum_{l\in I}a_{ik}m_{kl}b_{lj}=\sum_{l\in I}\sum_{k\in I}a_{ik}m_{kl}b_{lj}=((am)b)_{ij}.$ (14)
Since $I_{1}:=\{i\in I\mid\exists k\in I:\ a_{ik}\neq 0\}$ and $I_{2}:=\{j\in I\mid\exists l\in I:\ b_{lj}\neq 0\}$ are finite, then $\operatorname{supp}(amb)$ is finite as a subset of $I_{1}\times I_{2}$ by ˜14. Thus, $AMA\subseteq A$. The converse inclusion is obvious because $a_{i}=e_{i}\cdot e_{i}\cdot a_{i}$ for all $a_{i}\in e_{i}A$, so that $A=A\cdot A\cdot A\subseteq AMA$. ∎
Lemma 3.16.
Let * be an associative product on $A$. If * is $\mathrm{id}$-matching with $\cdot$, then $e_{i}*e_{j}\in e_{i}Ae_{j}$ for all $i,j\in I$.
Proof.
Since $e_{i}\cdot e_{i}=e_{i}$, then $e_{i}*e_{j}=(e_{i}\cdot e_{i})*e_{j}=e_{i}\cdot(e_{i}*e_{j})$. Similarly, $e_{j}\cdot e_{j}=e_{j}$ yields $e_{i}*e_{j}=e_{i}*(e_{j}\cdot e_{j})=(e_{i}*e_{j})\cdot e_{j}$. Hence $e_{i}*e_{j}=e_{i}\cdot(e_{i}*e_{j})\cdot e_{j}\in e_{i}Ae_{j}$. ∎
We thus define
$x=\prod_{i,j\in I}(e_{i}*e_{j})\in M.$ (15)
Lemma 3.17.
Let * be an associative product on $A$. If * is $\mathrm{id}$-matching with $\cdot$, then for all $a_{ij}\in e_{i}Ae_{j}$ and $a_{kl}\in e_{k}Ae_{l}$ we have $a_{ij}*a_{kl}=a_{ij}xa_{kl}$.
Proof.
By ˜15 and 14
$a_{ij}*a_{kl}$ $=(a_{ij}\cdot e_{j})*a_{kl}=a_{ij}\cdot(e_{j}*a_{kl})=a_{ij}\cdot(e_{j}*(e_{k}\cdot a_{kl}))$
$=a_{ij}\cdot((e_{j}*e_{k})\cdot a_{kl})=a_{ij}\cdot x_{jk}\cdot a_{kl}=a_{ij}xa_{kl}.\qed$
Proposition 3.18.
The $\mathrm{id}$-matching structures on $(A,\cdot)$ are exactly the products * on $A$ of the form
$a*b=axb$ (16)
for all $a,b\in A$.
Proof.
The fact that any $\mathrm{id}$-matching with $\cdot$ product * on $A$ has form ˜16 is a consequence of Lemma˜3.17. Conversely, let * be given by ˜16. Then
$(a*b)\cdot c=(axb)\cdot c=ax(b\cdot c)=a*(b\cdot c)$
and similarly
$a\cdot(b*c)=a\cdot(bxc)=(a\cdot b)xc=(a\cdot b)*c.\qed$
Lemma 3.19.
Let $\varphi:A\to A$ be a linear map. Then $\varphi\in\Gamma(A)$ if and only if there exists $m\in M$ such that
$\varphi(a)=ma=am$ (17)
for all $a\in A$.
Proof.
Let $\varphi\in\Gamma(A)$. Then $\varphi(e_{i})=\varphi(e_{i}\cdot e_{i})=e_{i}\varphi(e_{i})=\varphi(e_{i})e_{i}$, so that $\varphi(e_{i})\in e_{i}Ae_{i}$. Define $m=\prod_{i\in I}\varphi(e_{i})\in M$. For any $a_{ij}\in e_{i}Ae_{j}$ we have
$\varphi(a_{ij})=\varphi(e_{i}\cdot a_{ij})=\varphi(e_{i})a_{ij}=m_{ii}a_{ij}=ma_{ij}.$
On the other hand,
$\varphi(a_{ij})=\varphi(a_{ij}\cdot e_{j})=a_{ij}\varphi(e_{j})=a_{ij}m_{jj}=a_{ij}m.$
Thus, ˜17 holds for all $a\in A$.
Conversely, assume ˜17. Then
$\varphi(a)\cdot b=(ma)\cdot b=m(a\cdot b)=\varphi(a\cdot b)\quad\textup{and}\quad a\cdot\varphi(b)=a\cdot bm=(a\cdot b)m=\varphi(a\cdot b),$
for all $a,b\in A$, so $\varphi\in\Gamma(A)$. ∎
Lemma 3.20.
Let * be an associative product on $A$. If * is $(12)$-matching with $\cdot$, then $e_{i}*e_{j}=0$ for all $i\neq j$ and $e_{i}*e_{i}\in e_{i}Ae_{i}$ for all $i\in I$.
Proof.
Let $i\neq j$. Then $e_{i}*e_{j}=(e_{i}\cdot e_{i})*e_{j}=e_{i}*(e_{i}\cdot e_{j})=0$. It follows that $(e_{i}*e_{i})\cdot e_{j}=e_{i}\cdot(e_{i}*e_{j})=0$ and $e_{j}\cdot(e_{i}*e_{i})=(e_{j}*e_{i})\cdot e_{i}=0$, so that $e_{i}*e_{i}\in e_{i}Ae_{i}$. ∎
We thus define as in ˜15
$x=\prod_{i\in I}(e_{i}*e_{i})\in M.$ (18)
Lemma 3.21.
Let * be an associative product on $A$. If * is $(12)$-matching with $\cdot$, then
$xa=ax$ (19)
for all $a\in A$.
Proof.
We have $e_{k}\cdot(e_{i}*a_{ij})=(e_{k}*e_{i})\cdot a_{ij}=0$ for all $a_{ij}\in e_{i}Ae_{j}$ and $k\neq i$ by Lemma˜3.20, so that
$e_{i}*a_{ij}=e_{i}\cdot(e_{i}*a_{ij})=(e_{i}*e_{i})\cdot a_{ij}=x_{ii}a_{ij}=xa_{ij}.$ (20)
Similarly, $(a_{ij}*e_{j})\cdot e_{k}=a_{ij}\cdot(e_{j}*e_{k})=0$ for all $a_{ij}\in e_{i}Ae_{j}$ and $k\neq j$, so that
$a_{ij}*e_{j}=(a_{ij}*e_{j})\cdot e_{j}=a_{ij}\cdot(e_{j}*e_{j})=a_{ij}x_{jj}=a_{ij}x.$ (21)
But $e_{i}*a_{ij}=e_{i}*(a_{ij}\cdot e_{j})=(e_{i}\cdot a_{ij})*e_{j}=a_{ij}*e_{j}$, whence ˜19 by ˜20 and 21. ∎
Lemma 3.22.
Let * be an associative product on $A$. If * is $(12)$-matching with $\cdot$, then there exists $\varphi\in\Gamma(A)$ such that for all $a_{ij}\in e_{i}Ae_{j}$ and $a_{kl}\in e_{k}Ae_{l}$ we have
$a_{ij}*a_{kl}=\varphi(a_{ij}\cdot a_{kl}).$ (22)
Proof.
Let $\varphi(a)=xa$, where $x$ is given by ˜18. By Lemmas˜3.19 and 3.21 we have $\varphi\in\Gamma(A)$. We are left to prove ˜22. If $j\neq k$, then
$a_{ij}*a_{kl}=(a_{ij}\cdot e_{j})*a_{kl}=a_{ij}*(e_{j}\cdot a_{kl})=0=\varphi(a_{ij}\cdot a_{kl}).$
Now, by ˜19 and 21 we conclude that
$a_{ij}*a_{jk}=a_{ij}*(a_{jk}\cdot e_{k})=(a_{ij}\cdot a_{jk})*e_{k}=(a_{ij}\cdot a_{jk})x=\varphi(a_{ij}\cdot a_{jk}).\qed$
Proposition 3.23.
Let * be an associative product on $A$. Then the following are equivalent:
  1. (i)
    * is $(12)$-matching with $\cdot$;
  2. (ii)
    * is interchangeable with $\cdot$;
  3. (iii)
    * is totally compatible with $\cdot$;
  4. (iv)
    * is determined by some $\varphi\in\Gamma(A)$.
Proof.
[…]
4 $\sigma$-matching and interchangeable structures on free non-unital associative algebras
We begin with a general result.
Proposition 4.1.
Let $(A,\cdot)$ be an associative algebra without zero divisors. Then for an associative product * on $A$ the following are equivalent:
  1. (i)
    * is $(12)$-matching with $\cdot$;
  2. (ii)
    * is interchangeable with $\cdot$;
  3. (iii)
    * is totally compatible with $\cdot$.
Proof.
[…]
$a\cdot(a*(b\cdot c))=(a*a)\cdot(b\cdot c)=((a*a)\cdot b)\cdot c=(a\cdot(a*b))\cdot c=a\cdot((a*b)\cdot c).$
Since $(A,\cdot)$ has no zero divisors, $a\cdot(a*(b\cdot c))=a\cdot((a*b)\cdot c)$ implies $a*(b\cdot c)=(a*b)\cdot c$. Thus, * is totally compatible with $\cdot$ by Remark˜1.3˜(iv).
[…]
$a\cdot((a\cdot b)*c)=a*((a\cdot b)\cdot c)=a*(a\cdot(b\cdot c))=a\cdot(a*(b\cdot c)).$
As above, $a\cdot((a\cdot b)*c)=a\cdot(a*(b\cdot c))$ yields $(a\cdot b)*c=a*(b\cdot c)$, because $(A,\cdot)$ has no zero divisors. Thus, * is totally compatible with $\cdot$ by Remark˜1.3˜(ii).
[…]
Clearly, $\mathrm{id}$-matching structures on an algebra without zero divisors are not in general totally compatible: as a counterexample it is enough to take a non-commutative associative unital division algebra (see Corollary˜2.7).
Free algebras are a classical example of algebras with no zero divisors. In view of Proposition˜2.6, we will only be interested in the non-unital case.
4.1 The free non-unital associative algebra
Fix a set of variables $X$ with $|X|>1$. Let $X^{*}$ (resp. $X^{+}$) be the free monoid (resp. free semigroup) over $X$. The elements of $X^{*}$ (resp. $X^{+}$) are all the words (resp. non-empty words) over $X$, where the product of any two words $w_{1}$ and $w_{2}$ is their concatenation $w_{1}w_{2}$. Let $(K\langle X\rangle,\cdot)$ (resp. $(K\langle X\rangle^{+},\cdot)$) be the free associative (resp. free non-unital associative) $K$-algebra over $X$. Observe that $K\langle X\rangle$ (resp. $K\langle X\rangle^{+}$) is the semigroup $K$-algebra of $X^{*}$ (resp. $X^{+}$). Recall from [5, 6.1] that $K\langle X\rangle$ is a domain. Clearly, $K\langle X\rangle^{+}$ is an ideal in $K\langle X\rangle$ and
$K\langle X\rangle^{+}=\bigoplus_{x\in X}x\cdot K\langle X\rangle=\bigoplus_{x\in X}K\langle X\rangle\cdot x.$ (23)
Proposition 4.2.
The $\mathrm{id}$-matching structures on $(K\langle X\rangle^{+},\cdot)$ are in a one-to-one correspondence with maps $\star:X\times X\to K\langle X\rangle^{+}$,
$(x,y)\mapsto x\star y=\sum_{u\in X}u\cdot R_{x,y,u}=\sum_{v\in X}L_{x,y,v}\cdot v,$ (24)
where $R_{x,y,u},L_{x,y,v}\in K\langle X\rangle$ and
$\sum_{v\in X}L_{x,y,v}\cdot(v\star z)=\sum_{u\in X}(x\star u)\cdot R_{y,z,u}$ (25)
for all $x,y,z\in X$.
Proof.
Let * be an $\mathrm{id}$-matching structure on $(K\langle X\rangle^{+},\cdot)$.
Given $x,y\in X$, write $x*y=\sum_{u\in X}u\cdot R_{x,y,u}=\sum_{v\in X}L_{x,y,v}\cdot v$ according to ˜23. Then
$(x*y)*z=\sum_{v\in X}(L_{x,y,v}\cdot v)*z=\sum_{v\in X}L_{x,y,v}\cdot(v*z)$
and
$x*(y*z)=\sum_{u\in X}x*(u\cdot R_{y,z,u})=\sum_{u\in X}(x*u)\cdot R_{y,z,u}$
for all $x,y,z\in X$. Thus, defining $\star$ to be the restriction of * to $X\times X$, we get ˜24 and 25 by the associativity of *.
Conversely, let $\star:X\times X\to K\langle X\rangle^{+}$ satisfying ˜24 and 25. Given $a,b\in X^{+}$, there are unique $x,y\in X$ and $a_{1},b_{1}\in X^{*}$, such that $a=a_{1}x$ and $b=yb_{1}$. Then define
$a*b=a_{1}\cdot(x\star y)\cdot b_{1}.$ (26)
In particular, $x*y=x\star y$ for all $x,y\in X$. The product ˜26 uniquely extends to a bilinear product * on the whole $K\langle X\rangle^{+}$. Taking additionally $c\in X^{+}$, since $bc=yb_{1}c$, by ˜26 we have
$(a*b)\cdot c=(a_{1}\cdot(x\star y)\cdot b_{1})\cdot c=a_{1}\cdot(x\star y)\cdot(b_{1}c)=a*(bc).$ (27)
Moreover, if $b=b_{2}u$ and $c=zc_{1}$ with $u,z\in X$ and $b_{2},c_{1}\in X^{*}$, then
$(ab)*c=(ab_{2}u)*(zc_{1})=(ab_{2})\cdot(u\star z)\cdot c_{1}=a\cdot(b_{2}\cdot(u\star z)\cdot c_{1})=a\cdot(b*c).$ (28)
Thus, we are left to prove that * is associative. Observe from ˜27, 28 and 24 that ˜25 is equivalent to
$(x\star y)*z=x*(y\star z)$ (29)
for all $x,y,z\in X$. Let $a=a_{1}x$, $b=yb_{1}$ and $c=zc_{1}$ with $x,y,z\in X$ and $a_{1},b_{1},c_{1}\in X^{*}$. Then by ˜27 and 28 we have
$(a*b)*c=(a_{1}\cdot(x\star y)\cdot b_{1})*c=a_{1}\cdot(((x\star y)\cdot b_{1})*z)\cdot c_{1}.$ (30)
Case 1. $b_{1}$ is empty. Then using ˜26, 30, 29 and 28 we have
$(a*b)*c=a_{1}\cdot((x\star y)*z)\cdot c_{1}=a_{1}\cdot(x*(y\star z))\cdot c_{1}=(a_{1}x)*((y\star z)\cdot c_{1})=a*(b*c).$
Case 2. $b_{1}=b_{2}u$ for some $u\in X$ and $b_{2}\in X^{*}$. Then using ˜26, 30, 27 and 28 we have
$(a*b)*c$ $=a_{1}\cdot(((x\star y)\cdot b_{2}u)*z)\cdot c_{1}=a_{1}\cdot((x\star y)\cdot b_{2}\cdot(u*z))\cdot c_{1}$
$=(a_{1}\cdot(x\star y))\cdot(b_{2}\cdot(u\star z)\cdot c_{1})=(a*y)\cdot(b_{1}*c)$
$=a*(y\cdot(b_{1}*c))=a*((yb_{1})*c)=a*(b*c).\qed$
Remark 4.3.
We are unable to describe the maps $\star:X\times X\to K\langle X\rangle^{+}$ satisfying ˜24 and 25 in simpler terms, so let us just give some natural examples. Any associative product $\star:X\times X\to X$, i.e. a semigroup structure on $X$, is of this form. On the other hand, any mutation $x\star y=x\cdot_{p}y$, where $p\in K\langle X\rangle$, is also of this form. We suspect that there also exist examples of more complicated maps $\star$.
Lemma 4.4.
Let * be a totally compatible structure on $(K\langle X\rangle^{+},\cdot)$. Then there exists $\lambda\in K$ such that $x*y=\lambda(x\cdot y)$ for all $x,y\in X$.
Proof.
For all $x\neq y$ in $X$ we have
$x\cdot(x*y)=(x*x)\cdot y,$ (31)
so $x*y\in K\langle X\rangle\cdot y$ and $x*x\in x\cdot K\langle X\rangle$. Similarly,
$x\cdot(y*y)=(x*y)\cdot y$ (32)
implies $x*y\in x\cdot K\langle X\rangle$ and $x*x\in K\langle X\rangle\cdot x$. Thus, there exist $p_{x,y}\in K\langle X\rangle$, $x\neq y$, such that
$x*y=x\cdot p_{x,y}\cdot y,$ (33)
and $p_{x},q_{x}\in K\langle X\rangle$ such that
$x*x=p_{x}\cdot x=x\cdot q_{x}.$ (34)
It follows from $(x*x)\cdot x=x\cdot(x*x)$ that $p_{x}=q_{x}$, which commutes with $x$ by ˜34. By Bergman’s centralizer theorem [2] (or directly by induction on the degree of $p_{x}$) one concludes that $p_{x}\in K\langle x\rangle$. Now, ˜31 gives
$x\cdot p_{x,y}=q_{x}$ (35)
and ˜32 gives
$p_{x,y}\cdot y=p_{y},$ (36)
so that $p_{x,y}\in K\langle x\rangle\cap K\langle y\rangle=K$. It also follows from ˜35 that $p_{x,y}=p_{x,v}$ for all $y,v\neq x$. Similarly, ˜36 gives $p_{x,y}=p_{u,y}$ for all $x,u\neq y$. Let $u\neq v$ and $x\neq y$. If $u\neq y$, then $p_{u,v}=p_{u,y}=p_{x,y}$. If $v\neq x$, then $p_{u,v}=p_{x,v}=p_{x,y}$. If $(u,v)=(y,x)$, then $(x*y)\cdot x=x\cdot(u*v)$ yields $p_{x,y}\cdot y=y\cdot p_{u,v}$, whence $p_{x,y}=p_{u,v}$ because $p_{x,y},p_{u,v}\in K$. Thus, $p_{x,y}$ is a scalar $\lambda\in K$ that does not depend on $x$ and $y$, so the result follows by ˜33, 35 and 34. ∎
The result of the next corollary is probably well-known, but we couldn’t find a reference.
Corollary 4.5.
We have $\Gamma(K\langle X\rangle^{+})=K$.
Proof.
Let $\varphi\in\Gamma(K\langle X\rangle^{+})$ and $_{\varphi}$ be the corresponding totally compatible product as in Lemma˜2.3. Then there exists $\lambda\in K$ such that $\varphi(x\cdot y)=x*_{\varphi}y=\lambda(x\cdot y)$ for all $x,y\in X$ by Lemma˜$4.$ On the other hand, $\varphi(x\cdot y)=\varphi(x)\cdot y$, whence $\varphi(x)=\lambda x$ for all $x\in X$. Let now $a=xa_{1}\in X^{+}$, where $x\in X$ and $a_{1}\in X^{+}$. Then $\varphi(a)=\varphi(xa_{1})=\varphi(x)\cdot a_{1}=\lambda x\cdot a_{1}=\lambda a$. By linearity, $\varphi=\lambda\mathrm{id}$. ∎
Proposition 4.6.
Let * be an associative product on $K\langle X\rangle^{+}$. Then the following are equivalent:
  1. (i)
    * is $(12)$-matching with $\cdot$;
  2. (ii)
    * is interchangeable with $\cdot$;
  3. (iii)
    * is totally compatible with $\cdot$;
  4. (iv)
    * is determined by some $\varphi\in\Gamma(K\langle X\rangle^{+})$.
Proof.
[…]
$a*b$ $=a_{1}x*yb_{1}$
$=a_{1}\cdot(x*y)\cdot b_{1}$
$=a_{1}\cdot\lambda(x\cdot y)\cdot b_{1}$
$=\lambda(a_{1}x\cdot yb_{1})$
$=\lambda(a\cdot b)$
$=\varphi(a\cdot b).\qed$
4.2 The free non-unital commutative associative algebra
Fix a non-empty set of variables $X$. Let $(K[X],\cdot)$ (resp. $(K[X]^{+},\cdot)$) be the free commutative (resp. free non-unital commutative) associative $K$-algebra over $X$. Observe that $K[X]$ is the algebra of polynomials in $x\in X$ over $K$ and $K[X]^{+}$ is its ideal of polynomials with zero constant term. It is clear that $K[X]$ is an integral domain and $K[X]^{+}=\sum_{x\in X}xK[X]$ (the sum is not direct for $|X|>1$).
Lemma 4.7.
Let $\varphi:K[X]^{+}\to K[X]^{+}$ be a linear map. Then $\varphi\in\Gamma(K[X]^{+})$ if and only if there exists $p\in K[X]$ such that
$\varphi(a)=pa$ (37)
for all $a\in K[X]^{+}$.
Proof.
It is obvious that a map of the form ˜37 belongs to $\Gamma(K[X]^{+})$. Conversely, let $\varphi\in\Gamma(K[X]^{+})$.
Case 1. $X=\{x\}$. Then defining $p:=\frac{\varphi(x)}{x}\in K[X]$ we have
$\varphi(x^{n})=x^{n-1}\varphi(x)=x^{n-1}\cdot xp=x^{n}p$
for all $n\in\mathbb{Z}_{+}$, whence ˜37 by linearity.
Case 2. $|X|>1$. Then choosing $x\neq y$ in $X$ we have
$x\varphi(y)=\varphi(xy)=y\varphi(x).$ (38)
Since $x\neq y$, it follows that $\varphi(x)$ is divisible by $x$ (in $K[X]$), i.e. there exists $p_{x}\in K[X]$ such that $\varphi(x)=xp_{x}$. Then ˜38 gives $xyp_{y}=xyp_{x}$, so $p_{x}=p_{y}$, which will be denoted by $p$. We thus obtain ˜37 for $a$ being a variable. Now, for any $x\in X$ and $a_{1}\in K[X]^{+}$ we have $\varphi(xa_{1})=\varphi(x)a_{1}=pxa_{1}$, so ˜37 holds for monomials of higher degree and thus for all $a\in K[X]^{+}$ by linearity. ∎
Lemma 4.8.
Let * be an $\mathrm{id}$-matching structure on $(K[X]^{+},\cdot)$. Then for all $x,y\in X$ and $a_{1},b_{1}\in K[X]$ we have
$(xa_{1})*(yb_{1})=(x*y)a_{1}b_{1}.$
Proof.
We have $(xa_{1})*(yb_{1})=(a_{1}x)*(yb_{1})=a_{1}\cdot(x*(yb_{1}))=a_{1}\cdot(x*y)\cdot b_{1}=(x*y)a_{1}b_{1}$. ∎
Proposition 4.9.
Let $X=\{x\}$ and * be an associative product on $K[X]^{+}$. Then the following are equivalent:
  1. (i)
    * is $\mathrm{id}$-matching with $\cdot$;
  2. (ii)
    * is $(12)$-matching with $\cdot$;
  3. (iii)
    * is interchangeable with $\cdot$;
  4. (iv)
    * is totally compatible with $\cdot$;
  5. (v)
    there exists $p\in K[X]^{+}$ such that
    $a*b=\frac{a\cdot b}{x^{2}}\cdot p$ (39)
    for all $a,b\in K[X]^{+}$.
Proof.
[…]
$x^{m}*x^{n}=(x*x)x^{m+n-2}=\frac{x^{m}\cdot x^{n}}{x^{2}}(x*x).$
Denoting p:=xxp:=x*x, we get ˜39 by bilinearity of * and $\cdot$.
[…]
$(a*b)\cdot c=a*(b\cdot c)=(a\cdot b)*c=a\cdot(b*c)=\frac{a\cdot b\cdot c}{x^{2}}\cdot p$
for all $a,b,c\in K[X]^{+}$.
[…]
Remark 4.10.
Observe by Lemma˜4.7 that the product ˜39 from Proposition˜4.9˜(v) is determined by some $\varphi\in\Gamma(K[X]^{+})$ if and only if $\deg(p)\geq 2$.
Lemma 4.11.
Let $|X|>1$ and * be an $\mathrm{id}$-matching structure on $(K[X]^{+},\cdot)$. Then there exists $\varphi\in\Gamma(K[X]^{+})$ such that for all $x,y\in X$:
$x*y=\varphi(x\cdot y).$ (40)
Proof.
For all $x\neq y$ in $X$ we have $x\cdot(y*y)=(xy)*y=y\cdot(x*y)$. It follows that xyx*y is divisible by $x$ and yyy*y is divisible by $y$ (in $K[X]$). Similarly, changing the brackets in x(yx)=x(xy)x*(yx)=x*(xy), we get
$(x*y)\cdot x=(x*x)\cdot y,$ (41)
which shows that xyx*y is divisible by $y$ (in $K[X]$). Thus, for all $x\in X$ there is $p_{x}\in K[X]$ and for all $\in X, xyx\neq y$, there is $p_{x,y}\in K[X]$ such that
$x*x=xp_{x}\text{ and }x*y=xyp_{x,y}.$
Choosing another pair $u\neq v$ in $X$, we have
$(xu)*(yv)=x\cdot(u*(yv))=x\cdot(u*v)\cdot y=xyuvp_{u,v}$
and similarly
$(xu)*(yv)=u\cdot(x*y)\cdot v=xyuvp_{x,y}.$
Thus,
$p_{x,y}=p_{u,v}=:p,$
and ˜40 holds for $x\neq y$ and $\varphi(x)=px$. Moreover, ˜41 shows that $xyp_{x}=x^{2}yp$, whence
$p_{x}=xp,$
and ˜40 holds for $x=y$ and the same $\varphi$ too. ∎
Proposition 4.12.
Let $|X|>1$ and * be an associative product on $K[X]^{+}$. Then the following are equivalent:
  1. (i)
    * is $\mathrm{id}$-matching with $\cdot$;
  2. (ii)
    * is $(12)$-matching with $\cdot$;
  3. (iii)
    * is interchangeable with $\cdot$;
  4. (iv)
    * is totally compatible with $\cdot$;
  5. (v)
    * is determined by some $\varphi\in\Gamma(K[X]^{+})$.
Proof.
[…]
$(xa_{1})*(yb_{1})=(x*y)a_{1}b_{1}=\varphi(x\cdot y)\cdot a_{1}b_{1}=\varphi(xa_{1}\cdot yb_{1})$
for all $x,y\in X$ and $a_{1},b_{1}\in K[X]$. Then ˜(v) follows by bilinearity.
[…]
The remaining equivalences are Proposition˜4.1. ∎
Acknowledgements
This work was partially supported by CMUP, member of LASI, which is financed by national funds through FCT — Fundação para a Ciência e a Tecnologia, I.P., under the project with reference UIDB/00144/2020. The author is grateful to Ivan Kaygorodov for useful comments on associative algebras and to the referees whose suggestions helped to correct several typos and improved the readability of the paper.